Using the Substitution method, how would you solve x-y=12 and x+y=20?
Answer by TimothyLamb(4379)

x-y=12
x+y=20
---
put the system of linear equations into standard form
---
x-y=12
x+y=20
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 16
y = 4
---
now by substitution:
---
y = 20 - x
x - y = 12
x - (20 - x) = 12
x - 20 + x = 12
2x = 12 + 20
x = (12 + 20)/2
x = 16
y = 20 - 16
y = 4
---
solution by substitution:
x = 16
y = 4
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---